According to indexmundi.com China produces the largest amount of ammonia with 45,520 in thousand metric ton. The profitability of the ammonia production is around 90% of the total ammonia production cost which is about Php 10,000 per ton.
Refer to the balanced chemical equation: N2(g) + 3H2(g) ➡️ 2NH3(g)

How much is the production cost for this amount of ammonia?
1
Expert's answer
2020-11-25T02:18:32-0500

Solution:

The production cost of one ton of ammonia = 10000/1.9 = 5263.2 php per ton.

The production cost of the largest amount of ammonia = 5263.2×45,520,000 = 239,580,864,000 php.

Answer: The production cost of the largest amount of ammonia is 239,580,864,000 php.